OpenStudy (anonymous):

Please help ASAP!! Aimee and Ben are purchasing a condominium and are financing $365,000. The mortgage is a 15-year 3/1 ARM at 6.25% with a cap structure of 1.5/8. What will their payments be for the first 3 years?

OpenStudy (anonymous):

These are the answer choices.. $3,448.13 $3,435.66 $3,129.59 $2,247.37

OpenStudy (amistre64):

it will be the same payment structure for a fixed mortgage since the rates dont change till 3 years afterwards

OpenStudy (anonymous):

??

OpenStudy (amistre64):

you keep asking the same types of question over and over again but continue to display a total ignorance of any of the procedures for solutions that we have presented to you. It is hard to develop a solution process that you can learn by if you cannot retain any of the basic fundamental processes.

OpenStudy (amistre64):

there is a formula presented in your material that defines how to calcuate a fixed mortgage payment. Please review it.

OpenStudy (anonymous):

U keep attempting to explain the "same" process but every explanation seems to be different. I would like to learn it but it's hard when no one will explain the steps or the formula

OpenStudy (amistre64):

the process to develop the formula can be tedious; but the formula for a Payment is defined as: Beginning Balance, times rate/12, times (1+rate/12) raised to an expoent that is equal to the number of periods in the term of the loan:\[\Large B~(\frac{.0625}{12})(1+\frac{.0625}{12})^{(12*15)}\] this is then divided by 1 less then that last term: \[\Large \frac{(....)}{(1+\frac{.0625}{12})^{(12*15)}-1}\]

OpenStudy (amistre64):

or at least thats how they seem to present it in the texts that others have been using on here. I find this setup so be less intuitive tho.

OpenStudy (amistre64):

the way I set it up is from calculating our the Payment from the general setup .... which gets me this: http://www.wolframalpha.com/input/?i=365000%281%2B.0625%2F12%29%5E%2812*15%29*%281-%281%2B.0625%2F12%29%29%2F%281-%281%2B.0625%2F12%29%5E%2812*15%29%29

OpenStudy (anonymous):

You may have put numbers in that paragraph but it's saying "math processing error"

OpenStudy (amistre64):

i loathe that error, it means your web browser is not reading the latex code properly. Try hitting f5 to refresh your browser, if it still errors out, then its an issue with the browser itself not being compatible with this site.

OpenStudy (anonymous):

I'm on my phone

OpenStudy (amistre64):